# Proactive Model Adaptation Against Concept Drift for Online Time Series Forecasting

### v1

@@ 5.2.2. Efficiency
As depicted in Fig. 5, we compare the GPU memory occupation and inference latency of different model adaptation methods on Traffic, the largest time series dataset. We record the online latency over the four steps introduced in Sec. 2, including training on the latest training sample and inference on the test samples by adapted models. Here, SOLID only fine-tunes the final layer rather than the whole model, reducing the gradient computation cost. More specifically, SOLID computes gradients on several selected training samples iteratively instead of concurrently, so as to save GPU memory. By contrast, Proceed only uses the latest training sample and thereby achieves the lowest latency. Moreover, Proceed is faster and more lightweight than the online ensembling method OneNet.
@@ 6. Conclusion
In this work, we highlight that online time series forecasting inherently has a temporal gap between each test sample and available training data, where concept drift may well occur. Through empirical study, we found that this gap hinders the effectiveness of existing model adaptation methods which passively rely on the forecasting feedback on recent data. To address this issue, we propose a novel online model adaptation framework named Proceed, which proactively adapts the forecast model to the test concept before forecasting each test sample. Specifically, Proceed first fine-tunes the model on the latest acquired training sample, then extracts latent features from time series data to estimate the undergoing concept drift, and efficiently maps the estimated drift into parameter adjustments that are tailored for the test sample. Furthermore, we synthesize diverse concept drift and optimize Proceed with generalization capacity of mapping concept drift to beneficial parameter adjustments. Extensive experiments on five real-world time series datasets demonstrate that our proposed Proceed remarkably reduces the forecast errors of various forecast models and surpasses the state-of-the-art online learning methods. Our proactive model adaptation method provides a new direction in addressing continuous concept drift for online systems, which we wish to help any other prediction tasks with a feedback delay issue.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China (2022YFE0200500), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and SJTU Global Strategic Partnership Fund (2021 SJTU-HKUST).
@@ Appendix A Theoretical Analysis
In practice, we initialize the parameters ${\mathbf{W}}_{2}^{(\ell)}$ as an all-zero matrix, and the adaptation coefficients stem from zeros. Thus, we have $\mathcal{A}(\boldsymbol{\theta};\boldsymbol{\phi}_{0})=\boldsymbol{\theta}$, where $\mathcal{A}$ denotes the model adpater and $\boldsymbol{\phi}_{0}$ denotes the initialized parameters of $\mathcal{A}$. Given randomly shuffled historical data, we train the adapter parameters into $\boldsymbol{\phi}$ that is close to an optimum $\boldsymbol{\phi}^{*}$ and approximates the transition probability $P(\boldsymbol{\theta}_{t}\mid\boldsymbol{\theta}_{t^{\prime}}),\mathbf{X}_{t^{\prime}},\mathbf{X}_{t}$. Let $\hat{\boldsymbol{\theta}}_{t}$ denote $\mathcal{A}(\boldsymbol{\theta}_{t-H};\boldsymbol{\phi})$, $\boldsymbol{\theta}_{t}^{*}$ denote the optimal model parameters for forecasting $\mathcal{X}_{t}$, and $\boldsymbol{\theta}_{t}^{*}=\mathcal{A}(\boldsymbol{\theta}_{t-H};\boldsymbol{\phi}^{*})$. The adapted model parameters $\hat{\boldsymbol{\theta}}_{t}$ generated by the well-learned model adapter are expected to get closer to the optimal model parameters $\boldsymbol{\theta}_{t}^{*}$ than those generated by $\mathcal{A}(\cdot;\boldsymbol{\phi}_{0})$. Formally, we have
@@
(10) $\|\mathcal{F}(\mathbf{X}_{t};\hat{\boldsymbol{\theta}}_{t})-\mathcal{F}(\mathbf{X}_{t};\boldsymbol{\theta}^{*}_{t})\|$ $<\mathcal{L}_{upper}\|\hat{\boldsymbol{\theta}}_{t}-\boldsymbol{\theta}^{*}_{t}\|.$
Note that $\mathcal{F}$ is usually a neural network as a continuous function of $\boldsymbol{\theta}$. We define $\mathcal{S}(\boldsymbol{\theta}^{*}_{t},\Delta)$ as a sphere centering at $\boldsymbol{\theta}^{*}_{t}$ with a radius $\Delta\in\mathbb{R}^{+}$. When $\Delta$ approaches 0, due to the continuity of $\mathcal{F}$, the upper bound and lower bound of Lipschitz constant within $\mathcal{S}(\boldsymbol{\theta}^{*}_{t},\Delta)$ will become closer and finally identical, i.e., $\lim_{\Delta\to 0^{+}}\mathcal{L}_{upper}/\mathcal{L}_{lower}=1$. Moreover, we have known that $\|\boldsymbol{\theta}_{t-H}-\boldsymbol{\theta}^{*}_{t}\|/\|\hat{\boldsymbol{\theta}}_{t}-\boldsymbol{\theta}^{*}_{t}\|>1$ in Eq. (7). Therefore, there always exists a constant $\Delta>0$ such that $\mathcal{L}_{upper}/\mathcal{L}_{lower}<\|\boldsymbol{\theta}_{t-H}-\boldsymbol{\theta}^{*}_{t}\|/\|\hat{\boldsymbol{\theta}}_{t}-\boldsymbol{\theta}^{*}_{t}\|$, where $\boldsymbol{\theta}_{t-H},\hat{\boldsymbol{\theta}}_{t}\in\mathcal{S}(\boldsymbol{\theta}^{*}_{t},\Delta)$. Thus it is always possible to satisfy the following inequality:
@@ B.2. Data Adaptation
Apart from model adaptation, data adaptation is also a mainstream approach to the concept drift in time series forecasting, which is orthogonal to model adaptation. The goal of data adaptation is to normalize historical training data and future test data into a common data distribution (Kim et al., 2022; Fan et al., 2023; Liu et al., 2023), reducing overfitting risks. Among them, RevIN (Kim et al., 2022) is the most popular method that applies instance normalization to each time series lookback window and restores the statistics to the corresponding predictions, making each sample follow a similar distribution. Such normalization-based data adaptation methods mainly focus on the statistical changes in the mean and deviation of time series, while they overlook the distribution shifts in more complex temporal dependencies between time steps and spatial dependencies between variates (Zhao and Shen, 2024). Meanwhile, the forecast model may underfit the normalized time series, as the removed statistics may also serve as informative signals for prediction (Liu et al., 2022).
@@ C.1. Datasets
We list detailed descriptions of datasets used for experiments as follows.
@@ C.2. Evaluation Settings
To the best of our knowledge, there is no open-source evaluation framework that considers the temporal gap in online learning for time series forecasting. Though Zhang et al. (Zhang et al., 2023) noticed the feedback delay, they sidestep this issue by performing one prediction task every $H$ time steps and performing the next task until the ground truth is available. Concretely, they use $\mathbf{X}_{t-L:t}$ to forecast $\mathbf{X}_{t+1:t+H}$ and then use $\bf{X}_{t-L+H:t+H}$ to forecast $\bf{X}_{t+H+1:t+2H}$. In contrast, the common practice is to perform forecasting at each time step between $t$ and $t+H$, continually updating the previous predictions of $\mathbf{X}_{t-L+H:t+H}$. The reason is that short-term forecasting is typically easier than long-term forecasting. Therefore, in this work, we implement a more realistic evaluation framework for online learning where we perform forecasting at each time step, which remains consistent with the traditional evaluation setting.
@@ C.3. Implementation Details
Appendix D Further Experimental Results
D.1. Performance Gap in terms of MAE
@@
Method Dataset $H$ ETTh2 ETTm1 Weather ECL Traffic

### v2

@@ 5.2.2. Efficiency
Figure 5. Efficiency comparison on the Traffic dataset ($H=24$). The horizontal axis is the online forecasting latency (millisecond) between updating the model and obtaining online predictions. The vertical axis is the average MSE on test data. The size of each circle represents the peak amount of GPU memory occupation (GB).
As depicted in Fig. 5, we compare the GPU memory occupation and inference latency of different model adaptation methods on Traffic, the largest time series dataset. We record the online latency, including feedback-based model adaptation and any other additional steps. Here, SOLID only fine-tunes the final layer rather than the whole model, reducing the gradient computation cost. More specifically, SOLID and SOLID++ computes gradients on several selected training samples iteratively instead of concurrently, so as to save GPU memory. By contrast, Proceed only uses the latest training sample and thereby achieves the lowest latency. Moreover, Proceed is faster and more lightweight than the online ensembling method OneNet.
5.3. Visualization of the Representation Space
(a) Concepts
(b) Concept drifts
Figure 6. Visualization of concept representations and concept drift representations on ETTm1 ($H=24$).
To verify our assumption that online data have OOD concepts, we adopt t-SNE to visualize the representations of concepts $\mathbf{c}_{t}$ and concept drifts $\boldsymbol{\delta}_{t-H\to t}$ on the historical training data and online data. As shown in Fig. 6(a), there are a great number of OOD concepts on the online data, which are distinct from any historical concept. By contrast, as shown in Fig. 6(b), we encounter much fewer OOD concept drifts on the online data. These results support our intuition illustrated in Fig. 3. Hence, it is desirable for the adaptation generator to generate adaptations based on concept drifts instead of concepts.
@@ 6. Conclusion
In this work, we highlight that online time series forecasting inherently has a temporal gap between each test sample and available training data, where concept drift may well occur. Through empirical study, we found that this gap hinders the effectiveness of existing model adaptation methods which passively rely on the feedback in forecasting recent data. To address this issue, we propose a novel online model adaptation framework named Proceed, which proactively adapts the forecast model to the test concept before forecasting each test sample. Specifically, Proceed first fine-tunes the model on the latest acquired training sample, then extracts latent features from time series data to estimate the undergoing concept drift, and efficiently maps the estimated drift into parameter adjustments that are tailored for the test sample. Furthermore, we synthesize diverse concept drift and optimize Proceed with generalization capacity of mapping concept drift to beneficial parameter adjustments. Extensive experiments on five real-world time series datasets demonstrate that our proposed Proceed remarkably reduces the forecast errors of various forecast models and surpasses the state-of-the-art online learning methods. Our proactive model adaptation method provides a new direction in addressing continuous concept drift for online systems, which we wish to help any other prediction tasks with a feedback delay issue.
@@ Acknowledgements.
This work is supported by the National Key Research and Development Program of China (2022YFE0200500), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and SJTU Global Strategic Partnership Fund (2021 SJTU-HKUST). We would like to thank Shaofeng Cai for his valuable advice on the preliminary version of the paper.
@@ Appendix A Theoretical Analysis
In practice, we initialize the parameters ${\mathbf{W}}_{2}^{(\ell)}$ as an all-zero matrix, and the adaptation coefficients stem from zeros. Thus, we have $\mathcal{A}(\boldsymbol{\theta};\boldsymbol{\phi}_{0})=\boldsymbol{\theta}$, where $\mathcal{A}$ denotes the model adpater and $\boldsymbol{\phi}_{0}$ denotes the initialized parameters of $\mathcal{A}$. Given randomly shuffled historical data, we train the adapter parameters into $\boldsymbol{\phi}$ that is close to an optimum $\boldsymbol{\phi}^{*}$ and approximates the transition probability $P(\boldsymbol{\theta}_{t}\mid\boldsymbol{\theta}_{s},\mathbf{X}_{s},\mathbf{X}_{t})$. Let $\hat{\boldsymbol{\theta}}_{t}$ denote $\mathcal{A}(\boldsymbol{\theta}_{t-H};\boldsymbol{\phi})$, $\boldsymbol{\theta}_{t}^{*}$ denote the optimal model parameters for forecasting $\mathbf{Y}_{t}$, and $\boldsymbol{\theta}_{t}^{*}=\mathcal{A}(\boldsymbol{\theta}_{t-H};\boldsymbol{\phi}^{*})$. The adapted model parameters $\hat{\boldsymbol{\theta}}_{t}$ generated by the well-learned model adapter are expected to get closer to the optimal model parameters $\boldsymbol{\theta}_{t}^{*}$ than those generated by $\mathcal{A}(\cdot;\boldsymbol{\phi}_{0})$. Formally, we have
@@
$\|\mathcal{F}(\mathbf{X}_{t};\hat{\boldsymbol{\theta}}_{t})-\mathcal{F}(\mathbf{X}_{t};\boldsymbol{\theta}^{*}_{t})\|$ $<\mathcal{L}_{upper}\|\hat{\boldsymbol{\theta}}_{t}-\boldsymbol{\theta}^{*}_{t}\|.$
Note that $\mathcal{F}$ is usually a neural network as a continuous function of $\boldsymbol{\theta}$. We define $\mathcal{S}(\boldsymbol{\theta}^{*}_{t},\Delta)$ as a sphere centering at $\boldsymbol{\theta}^{*}_{t}$ with a radius $\Delta\in\mathbb{R}^{+}$. When $\Delta$ approaches 0, due to the continuity of $\mathcal{F}$, the upper bound and lower bound of Lipschitz constant within $\mathcal{S}(\boldsymbol{\theta}^{*}_{t},\Delta)$ will become closer and finally identical, i.e., $\lim_{\Delta\to 0^{+}}\mathcal{L}_{upper}/\mathcal{L}_{lower}=1$. Moreover, we have known that $\|\boldsymbol{\theta}_{t-H}-\boldsymbol{\theta}^{*}_{t}\|/\|\hat{\boldsymbol{\theta}}_{t}-\boldsymbol{\theta}^{*}_{t}\|>1$ in Eq. (7). Therefore, there exists a constant $\Delta>0$ such that $\mathcal{L}_{upper}/\mathcal{L}_{lower}<\|\boldsymbol{\theta}_{t-H}-\boldsymbol{\theta}^{*}_{t}\|/\|\hat{\boldsymbol{\theta}}_{t}-\boldsymbol{\theta}^{*}_{t}\|$, where $\boldsymbol{\theta}_{t-H},\hat{\boldsymbol{\theta}}_{t}\in\mathcal{S}(\boldsymbol{\theta}^{*}_{t},\Delta)$. Thus it is possible to satisfy the following inequality:
@@ B.2. Data Adaptation
Apart from model adaptation, data adaptation is another mainstream approach to concept drift in time series forecasting, which is orthogonal to model adaptation. The goal of data adaptation is to normalize historical training data and future test data into a common data distribution (Kim et al., 2022; Fan et al., 2023; Liu et al., 2023), reducing overfitting risks. Among them, RevIN (Kim et al., 2022) is the most popular method that applies instance normalization to each time series lookback window and restores the statistics to the corresponding predictions, making each sample follow a similar distribution. Such normalization-based data adaptation methods mainly focus on the statistical changes in the mean and deviation of time series, while they overlook the distribution shifts in more complex temporal dependencies between time steps and spatial dependencies between variates (Zhao and Shen, 2024). Meanwhile, the forecast model may underfit the normalized time series, as the removed statistics can serve as informative signals for prediction (Liu et al., 2022).
@@ C.1. Datasets
We list descriptions of datasets used for experiments as follows.
@@ C.2. Evaluation Settings
To the best of our knowledge, there is no open-source evaluation framework that considers the temporal gap in online learning for time series forecasting. Though Zhang et al. (Zhang et al., 2023) noticed the feedback delay, they sidestep this issue by performing one prediction task every $H$ time steps and performing the next task until the ground truth is available. Concretely, they use $\mathbf{X}_{t-L:t}$ to forecast $\mathbf{X}_{t+1:t+H}$ and then use $\mathbf{X}_{t-L+H:t+H}$ to forecast $\mathbf{X}_{t+H+1:t+2H}$. In contrast, the common practice is to perform forecasting at each time step between $t$ and $t+H$, continually updating the previous predictions of $\mathbf{X}_{t-L+H:t+H}$. The reason is that short-term forecasting is typically easier than long-term forecasting. Therefore, in this work, we implement a more realistic evaluation framework for online learning where we perform forecasting at each time step, which remains consistent with the traditional evaluation setting.
@@ C.3. Implementation Details
Dataset $H$ ETTh2 ETTm1 Weather ECL Traffic
@@
Appendix D Pipeline of Proceed
Figure 7. Pipeline of Proceed at each online time $t$. CDE: concept drift estimation; PMA: proactive model adaptation.
Fig. 7 depicts the pipeline of Proceed during the online phase, which is a bit different from the steps introduced in Sec. 4.1. Given model parameters $\boldsymbol{\theta}_{t-H-1}$ that have been updated with previous training data $\mathcal{D}_{(t-1)-}$, our model adapter additionally estimates the concept drift between $\mathcal{D}_{(t-1)-}$ and $\mathcal{D}_{t-}$ and adapts the model to $\mathcal{D}_{t-}$, yielding forecast feedback for updates. The reason behind it is that we jointly train the forecast model and the model adapter (see Sec. 4.4), while our training algorithm does not enforce the forecast model to perform well alone. Consequently, the forecast model without $\Delta\boldsymbol{\theta}$ could forget some predictive skills, while the parameter adjustments $\Delta\boldsymbol{\theta}$ from the model adapter may incorporate the skills forgot by the model. Thus it is necessary for the inference process to keep consistent with the training phase, i.e., we always make predictions by integrating the forecast model and our model adapter as a whole. Nevertheless, in the feedback-based adaptation step, we only finetune the forecast model but keep the model adapter’s parameters frozen. Otherwise, the model adapter may overfit the patterns of the one-step concept drift between $\mathcal{D}_{(t-1)-}$ and $\mathcal{D}_{t-}$, losing generalization ability.
Appendix E Further Experimental Results
E.1. Performance Gap in terms of MAE
E.2. Performance with Partial Ground Truth
Table 8. MSE when using $\mathbf{Y}_{t-H}$ and partial ground-truth.
Method Dataset $H$ ETTh2 ETTm1 Weather
Model 24 48 96 24 48 96 24 48 96
PatchTST GD 1.833 3.144 5.429 0.455 0.591 0.673 0.734 0.977 1.261
Proceed 1.703 3.070 5.299 0.420 0.570 0.650 0.719 0.967 1.256
iTransfo. GD 2.463 3.759 6.152 0.459 0.602 0.692 0.851 1.131 1.406
Proceed 2.228 3.571 5.868 0.426 0.569 0.659 0.734 1.001 1.299
In this section, we introduce variants that update the model with feedback on $\mathbf{Y}_{t-H}$ and partial ground truth $\{\tilde{\mathbf{Y}}_{t-i}\}_{i=1}^{H-1}$, where ${\tilde{\mathbf{Y}}_{t-i}=[v_{t-i},\cdots,v_{t-1}]}$. Since each update involves $H$ training samples, the GPU memory overhead becomes about $H$ times than one involving $\mathbf{Y}_{t-H}$ only. We leave out results on ECL and Traffic datasets, where the GPU memory overhead exceeds the limit of our Nvidia 4090 GPUs (24GB). As shown in Table 8, Proceed consistently outperforms online gradient descent method under this setting. Compared with Table 3, using both $\mathbf{Y}_{t-H}$ and partial ground truth for feedback-based adaptation can improve the performance, while the improvement is insignificant. This indicates that samples with partial ground truth cannot reveal the concept of the test sample, necessitating proactive model adaptation. Though the partial ground truth can be beneficial to our method, the finetuning cost scales up with $H$ (e.g., up to either 96$\times$ latency or 96$\times$ GPU memory). Thus, it would be much more efficient if using $\mathbf{Y}_{t-H}$ only. Notably, our proposed method using $\mathbf{Y}_{t-H}$ only can still outperform GD using $H$ training samples in most cases. .
E.3. Variants of the Concept Encoder
In this section, we study different kinds of operations on multivariate time series using our concept encoder. Let $\mathbf{c}_{t,i}$ denote the latent concept feature vector of the $i$-th variate. Apart from the average introduced in Eq. (1), we implement two variants as follows:
  • linear transformation: we learn a linear layer that transforms the concatenation of all latent features, i.e., $c_{t}=W^{\top}[c_{t0},c_{t1},\cdots]$.
  • weighted sum: we learn an individual weight $w_{i}\in\mathbb{R}$ for each $i$-th variate, and $c_{t}=\sum_{i}w_{i}c_{ti}$.
In Table 9, we compare the performance of the variants in terms of the average MSE with $H\in$ 24,48,96. The results demonstrate that the average operation is simple yet effective. One possible reason is that all variates are typically assumed to have equal importance during both training and evaluation.
Table 9. Performance comparison between variants of concept encoders..
Model Method ETTh2 ETTm1 Weather ECL Traffic
PatchTST linear trans. 3.498 0.570 0.985 5.661 0.363
weighted sum 3.467 0.571 0.986 5.435 0.362
average 3.468 0.554 0.986 4.732 0.356
iTransfo. linear trans. 4.400 0.548 1.018 4.944 0.355
weighted sum 4.441 0.551 1.021 5.260 0.352
average 4.216 0.543 1.017 4.752 0.352